if one painter can paint a house in 5 days and another painter can paint the same house in 4 days, how long will it take them to paint the house together?
Answer by edjones(7569)

The 1st can do 1/5 the job in a day and the 2nd 1/4 the job in a day.
Together:
1/5 + 1/4 = 1/t
4/20 + 5/20 = 1/t
9/20=1/t
t=20/9 hrs.
.
Ed